OpenStudy (anonymous):

I need help with this integral: integral [e^-sqrt(x)]/[sqrt(x)] from 1 to infinity.

OpenStudy (anonymous):

u =

OpenStudy (anonymous):

see what you get when you make your -sqrt(x) your u

myininaya (myininaya):

\[\int\limits_{}^{} \frac{e^{-\sqrt{x}}}{\sqrt{x}} dx \] \[\text{Let } u=\sqrt{x} => du=\frac{1}{2 \sqrt{x}} dx => 2 du=\frac{1}{\sqrt{x}} dx\]

OpenStudy (anonymous):

I did do the u-sub, I just didn't know where to go from there. I really suck at improper integrals.

myininaya (myininaya):

\[\int\limits_{}^{}e^{-u} 2 du =2(-e^{-u})+C=-2e^{-u}+C =-2e^{-\sqrt{x}}+C\] so we have \[\int\limits_{1}^{\infty}\frac{e^{-\sqrt{x}}}{\sqrt{x}} dx=\lim_{b \rightarrow \infty}[-2e^{- \sqrt{b}}-(-2e^{- \sqrt{1}})]\]

myininaya (myininaya):

recall \[\lim_{b \rightarrow \infty}e^{-b}=0\]

OpenStudy (anonymous):

you don't have to recall simply move your e^-sq(x) to the bottom

OpenStudy (anonymous):

Okay thank you so much! That's very helpful :)

myininaya (myininaya):

well i would recall it is easy

OpenStudy (anonymous):

\[-2e^{-x}+c=\frac{-2}{e^\sqrt{x}}\]

myininaya (myininaya):

\[-2(0)+2\frac{1}{e}=\frac{2}{e}\]

OpenStudy (anonymous):

Oh wait, I just got confused. What happens to the sqrt(x) on the bottom after you do the u-sub?

myininaya (myininaya):

do you remember what was du?

OpenStudy (anonymous):

myin has it correct but if you didn't know that e^-b goes to 0, you can just simply do what i did and solve to see that to infinity goes to 0

OpenStudy (anonymous):

Oh right! Yes okay I see now. Sorry I forgot to move that over w/ the 2.

OpenStudy (anonymous):

how long is it going to say "website is restarting lol"

myininaya (myininaya):

its annoying

OpenStudy (anonymous):

Oh okay, my calc one teacher drilled it into my head, so I remember that negative infinity goes to zero when it's an exponent :) Thank tho!

OpenStudy (anonymous):

lol i know haha how do you know that e^-b goes to 0 ... do you read your calc book for fun lmfao =]

myininaya (myininaya):

oh wait i exited out of it and it didn't pop back up oh wait it just popped back up :(

OpenStudy (anonymous):

how long has that sign been up? I'm guessing longer than an hour?

myininaya (myininaya):

e^{x} goes to 0 as x goes to negative infinity so e^{-x} goes to 0 as x goes to infinity

OpenStudy (anonymous):

yes because \[x^{-\infty}=\frac{1}{x^\infty}\]

myininaya (myininaya):